OpenStudy (anonymous):

Evaluate x divided by 80 for x = 8/9. The choices are A.71 1/9 B.1/9 C.9/640 D.90

jhonyy9 (jhonyy9):

so what are given there above you can rewrite them so x 8 --- = ---- cross multiplie and will get 80 9 9x= 8*80 9x=640 x = 640/9 = so can you ending it now alone ?

OpenStudy (jdoe0001):

hmmm I understood \(\bf \cfrac{x}{80}\qquad x=\cfrac{8}{9}\implies \cfrac{\frac{8}{9}}{80}\implies \cfrac{\frac{8}{9}}{\frac{80}{1}} \\ \quad \\ \cfrac{\cancel{8}}{9}\cdot {\color{brown}{ \cfrac{1}{\cancel{80}} }}\implies ?\)

OpenStudy (jdoe0001):

which gives none of the choices provided maybe if you can provide us with a screenshot of the material

jhonyy9 (jhonyy9):

@jdoe0001 - sorry for your derange but than you see ,yes i know my wrong answer but how i have got there above x = 640 / 9 so from this result = 71 1/9 right ?

jhonyy9 (jhonyy9):

@Indiya - do you see it now what i have wrote above to @jdoe0001 ????

jhonyy9 (jhonyy9):

so from this result the choice A is right - ok ?